Institut National de la Recherche en Informatique et en Automatique

scilab−4.1.2

--> ↩ ↑

↓ rpertoire courant }

getcwd()
pwd()
cd()
chdir()
mkdir()
ls

+−∗/

\

(1/2), sqrt(2)ans = e π %e,%pi

%i i %eps

%eps
%inf

10^{300} %inf %T %F

(160))^{2}ans =

Inf

+*/^[](){}'&|~ ans palmodelica_{l}ibsscicos_{p}al_{l}ibshomeSCIHOMEP W DT MP DIRMSDOSSCIguilibsparselibxdesslibpercentlibpolylibintlibelemlibutillibstatslibalglibsigliboptlibautolibroblibsoundlibmetalibarmalibtkscilibtdcslibs2flibmtlblibusing18185elementsoutof5000000.and71variablesoutof9231

i, j

A A(i,j)

A(i,j)=

2 −1 (p, n)

.ˆ .ˆ

2ans = Aans = ^ˆ

2ans = 1× 1 ⇔ ⇎

A^(-1)

(− 1)ans =

length()

%T %F <><><=>=|∼

<><=>=<>|˜

length()

k^{ieme}

part(chane,k)

\begin{matrix}f : & A & ⟶ & B \\ x & ⟼ & y = f(x) \end{matrix}

f(x) = 1+

x^{2} 2endfunction

2′)//dfinitionenligne− − > f(2)ans =

f(x)=1+x^2 x^2 2ans =

if condition then action1
             else action2
end
for cpt=debut:pas:fin
    actions
end
while condition
    actions

end condition action∗ n! = n× (n−

1)!

1! = 1 factoriel(n) =

n!